# Deep multi-intentional inverse reinforcement learning for cognitive multi-function radar inverse cognition

### v1

@@ 4.1 Simulation Settings
, where $w_{1}$, $w_{2}$, $w_{3}$ are the positive weights that determine the importance of each quantity. In general, we encourage the radar to maximize its detection or tracking performance while lowering the transmission power. Depending on the specific application (e.g., Search, Track) or goals (e.g., energy minimization), different weights could be applied [23]. Table 2 summarizes typical choices of these parameters under specific applications.
@@
In this study, we consider a CFMR with a 2-layer hierarchy that can switch between 3 tasks. Task 1 is for target tracking, with corresponding weights set to $w_{1}=5$, $w_{2}=5$, $w_{3}=1$; Task 2 is for LPI, with corresponding weights set to $w_{1}=1$, $w_{2}=5$, $w_{3}=5$; Task 3 is for target classification, with corresponding weights set to $w_{1}=5$, $w_{2}=1$, $w_{3}=1$. For all the tasks, the radar interacts with the environment with random actions to collect 10,000 trajectories to estimate the transition matrix $P$, and the value iteration algorithm is applied to find a near-optimal policy.
@@ 4.2 Baseline Methods and Evaluation Metrics
  • Deep MaxEnt IRL[35]: An extension of MaxEnt IRL that uses deep neural networks to approximate the reward function.
@@
  • RNN: A widely used neural network architecture for sequence modeling. This model has been used in [18, 26, 19] for radar work modes (resource management) recognition.
@@
  • Action Prediction Error: This metric measures the deviation of the actions predicted by the inferred policy from the actions taken in the true trajectories. It is defined as follows:
@@
4.3 Result on single Intention IRL
@@
First, we conduct the experiment in an ideal situation, where the ESM can correctly extract all the actions of the radar. 1000 demonstration trajectories are used, and the training curves are presented in Fig6.
Figure 6: EVD of different methods at a varying number of training iterations.
As observed, the model employing a deep neural network shows superior performance, quickly reducing the EVD to a low value. In contrast, the algorithm with linear weights has notably higher EVD throughout the training process. This indicates that a neural network is necessary when inferring such complex reward structures. Compared with DMLIRL, DmaxEnt IRL, while still converging, does so more slowly. This indicates that directly backpropagating through the soft policy may be more efficient.
Next, we test the performance under different ratios of action extraction errors. It is assumed that each action is equally likely to be flipped to a wrong one due to non-ideal receiving of the ESM. We use the Action Error Ratio (AER) to represent the probability that any given action is misinterpreted. As observed in Fig. 7, the EVD increases as the Action Error Ratio increases for all methods, indicating a degradation in performance as the action extraction becomes noisier. The DMLIRL algorithm with $\beta=0.5$ shows the most robust performance, highlighting that we could tune the parameter $\beta$ in Maximum Likelihood IRL to cope with various non-ideal conditions.
@@
Lastly, we limited the number of available trajectories to investigate how different IRL algorithms perform with varying amounts of training data. The number of training trajectories ranged from 10 to 600, and we measured both the EVD and the Action Prediction Error (APE) to assess the effectiveness of each algorithm.
The results are presented in Fig. 8. Both EVD and APE decrease sharply as the number of trajectories increases. This is because both the estimation of the transition matrix and the diversity of the expert demonstrations can affect the performance. The more training trajectories available, the more accurate the transition matrix estimation becomes, and the more diverse the expert demonstrations are, leading to better generalization and reduced error in policy prediction. Noticeably, when the dataset contains fewer than 100 trajectories, the Maximum-Entropy-based method outperforms the Maximum-Likelihood-based method. This suggests that Maximum-Entropy methods are more effective in handling smaller datasets, as maximizing the entropy makes minimal commitments beyond the constraints of feature matching.
Figure 8: EVD of different methods at various action prediction error rates.
@@ 4.4 Results on multiple Intentions IRL
In this section, we assume the radar can perform different tasks in different trajectories. Thus, the dataset is generated by Task 1, 2, and 3, and the reconnaissance side has to not only estimate the reward functions but also cluster the trajectories. 1000 demonstration trajectories for each task are generated and mixed to evaluate the performance, and the results are summarized in Table 3.
@@
Method NMI ARI EVD APE(%)
DMLIRL - - 32.45 74.23
DmaxEnt IRL - - 24.75 69.47
DMLIRL+EM 0.99 0.99 3.45 5.20
DmaxEnt+EM 0.95 0.89 5.61 7.27
RNN 0.88 0.83 - -
K-means 0.50 0.37 - -
GMM 0.56 0.30 - -
For clustering, the DMLIRL+EM method demonstrates the highest clustering accuracy with both NMI and ARI values at 0.99, indicating nearly perfect alignment with the ground truth task labels. Similarly, DmaxEnt+EM also performs well, with an NMI of 0.95 and an ARI of 0.89. The RNN method, while only using actions as input, achieves reasonably good results because it uses labels to guide the learning process. In contrast, traditional clustering methods like K-means and GMM show significantly lower performance because they are purely unsupervised and do not utilize any sequential information.
For reward estimation, the methods with just a single model failed to converge as the EVD and APE are significantly higher than those of multiple intentional IRL models. These results show that datasets produced by multiple experts cannot be well explained with a single model.
@@ 5 Conclusion
The paper introduces a deep multi-intentional inverse reinforcement learning method for effectively understanding and analyzing cognitive multi-function radars (CMFR). By leveraging an EM algorithm with neural networks, the proposed method successfully clusters trajectories and estimates multiple reward functions in the presence of complex and multi-task radar settings. The simulation results showcase the robustness and superior performance of DMLIRL+EM in comparison to single expert IRL methods and traditional pattern recognition approaches.

### v2

@@ 4.1 Simulation Settings
, where $w_{1}$, $w_{2}$, $w_{3}$ are the positive weights that determine the importance of each quantity. In general, we encourage the radar to maximize its detection or tracking performance while lowering the transmission power. Depending on the specific application (e.g., Search, Track) or goals (e.g., energy minimization), different weights could be applied [23]. Table 2 summarizes typical choices of these parameters under specific goals.
@@
Spectrum sharing Small Small Large
In this study, we consider a CFMR with a 2-layer hierarchy that can switch between 3 tasks. Task 1 is for target searching, with corresponding weights set to $w_{1}=5$, $w_{2}=1$, $w_{3}=5$; Task 2 is for LPI, with corresponding weights set to $w_{1}=1$, $w_{2}=5$, $w_{3}=5$; Task 3 is for Spectrum sharing, with corresponding weights set to $w_{1}=1$, $w_{2}=1$, $w_{3}=5$. For all the tasks, the radar interacts with the environment with random actions to collect 10,000 trajectories to estimate the transition matrix $P$, and the value iteration algorithm is applied to find a near-optimal policy.
@@ 4.2 Baseline Methods and Evaluation Metrics
  • Deep MaxEnt IRL[35]: An extension of MaxEnt IRL that uses deep neural networks to approximate the reward function. This method has been applied to estimate rewards of CR in [36].
@@
  • RNN: A widely used neural network architecture for sequence modeling. This model has been used in [18, 26, 19] for radar work modes (resource management) recognition, and used in [26] for complex radar emitter behavior prediction.
@@
  • Action Prediction Error(APE): This metric measures the deviation of the actions predicted by the inferred policy from the actions taken in the true trajectories. It is defined as follows:
@@
4.3 Results on single Intention IRL
@@
First, we conduct the experiment in an ideal situation, where the ESM can correctly extract all the actions of the radar. 1000 demonstration trajectories are used, and the training curves are presented in Fig. 6.
Figure 6: EVD and APE of different methods at a varying number of training iterations.
As observed, the model employing a deep neural network shows superior performance, quickly reducing the EVD to a low value. In contrast, the algorithm with linear weights completely failed to converge. This indicates that a neural network is necessary when inferring such complex reward structures. DMLIRL and DmaxEnt IRL, both method have nearly the same convergence speed and can completely replicate the behavior of the expert.
Next, we test the performance under different ratios of action extraction errors. We only test for DMLIRL and DmaxEnt IRL as other methods failed to converge. It is assumed that each action is equally likely to be flipped to a wrong one due to non-ideal receiving of the ESM. We use the Action Error Ratio (AER) to represent the probability that any given action is misinterpreted. As observed in Fig. 7, the EVD only slightly increased as the Action Error Ratio increases, up to 0.3. This indicates that both DMLIRL and DmaxEnt IRL is not sensitive to noisy observations. DmaxEnt IRL slightly outperforms DMLIRL as maximizing the entropy makes minimal commitments and is the least wrong[aroraSurveyInverseReinforcement2021].
@@
Lastly, we limited the number of available trajectories to investigate how different IRL algorithms perform with varying amounts of training data. The number of training trajectories ranged from 10 to 1000, and we measured both the EVD and the Action Prediction Error (APE) to assess the effectiveness of each algorithm.
The results are presented in Fig. 8. Both EVD and APE decrease sharply as the number of trajectories increases. This is because both the estimation of the transition matrix and the diversity of the expert demonstrations can affect the performance. The more training trajectories available, the more accurate the transition matrix estimation becomes, and the more diverse the expert demonstrations are, leading to better generalization and reduced error in policy prediction. Noticeably, when the dataset contains fewer than 300 trajectories, the Maximum-Likelihood-based method outperforms the Maximum-Entropy-based method. This is because in ML-based IRL, the estimated transition matrix only affect the training through value iteration, whereas for Maxent-based IRL, both the value iteration and the policy propagation (see algorithm 2) can be affected.
Figure 8: EVD and APE of different methods under different number of training trajectories.
@@ 4.4 Results on multiple Intentions IRL
In this section, we assume the radar can perform different tasks in different trajectories. Thus, the dataset is generated by Task 1, 2, and 3, and the reconnaissance side has to not only estimate the reward functions but also cluster the trajectories. 1000 demonstration trajectories for each task are generated and mixed to evaluate the performance. The training curves for different IRL methods are presented in Fig. 9, and final testing results are summarized in Table 3.
Figure 9: Training curves for different IRL method on the dataset containing task 1,2 and 3.
@@
Method NMI ARI EVD APE
DMLIRL - - 26.21 0.48
DmaxEnt IRL - - 41.36 0.54
MI-DMLIRL 1.00 1.00 4.26 0.07
MI-DmaxEnt 0.73 0.57 18.97 0.17
RNN 1.00 1.00 - 0.14
K-means 0.57 0.47 - -
GMM 0.63 0.54 - -
For reward estimation, the methods with just a single model failed to converge as the EVD is significantly higher than those of multiple intentional IRL models. This is because different trajectories of the CMFR can come from completely different tasks, which can not be explained with a reward model. Compared with the Maxent-based IRL, ML-based IRL better suits EM based the training. As shown in Fig. 9, for MI-DMIRL both EVD and APE decreased significantly for each EM iteration and quickly converge to near optimal values. However, the training for MI-DmaxentIRL is not so stable, and it finally stuck in a local minimal.
For clustering, the MI-DMLIRL method demonstrates perfect clustering with both NMI and ARI values reach 1, which is the same as the supervised-trained RNN. In contrast, MI-DmaxEnt fails to make perfect alignment with the ground truth task labels as its learning stuck at local minimal. However, it still performances better than traditional clustering methods such as K-means and GMM because they only use the action information.
For actions prediction, the MI-DMLIRL method excels in predicting the correct actions, as evidenced by its low APE of 0.07. In contrast, MI-DmaxEnt shows a higher APE of 0.17 as its convergence to suboptimal solutions impairs its ability to consistently predict actions across all tasks. Interestingly, the RNN model, although not using the state-action relations, also achieves a low APE of 0.14. This again proofs that context information is important for complex radar emitter behavior prediction[apfeldValueMemoryMarkov2021].
@@ 5 Conclusion
The paper introduces a deep multi-intentional inverse reinforcement learning method for effectively understanding and analyzing cognitive multi-function radars. By leveraging an EM algorithm with neural networks, the proposed method successfully clusters trajectories and estimates multiple reward functions in the presence of complex and multi-task radar settings. The simulation results showcase the robustness and superior performance of DMLIRL+EM in comparison to single expert IRL methods and traditional pattern recognition approaches.